OpenStudy (anonymous):

Solve the equation by completing the square? 12x= -2x² + 8 A. x= -3 + √13, -3, - √13 B. x= 3 + √13, 3 -√13 C. x= 2,4 D. x= -2,-4

OpenStudy (saifoo.khan):

Learn here; http://www.youtube.com/watch?v=gzm-uhj06q8

OpenStudy (anonymous):

\[2x ^{2}+12x-8=2(x+3)^2-26=0 => (x+3)^2=13 => x=-3+\sqrt{13} , x=-3-\sqrt{13}\]

OpenStudy (anonymous):

So B?